I have such a hard time doing this problem:
Write Standard Equation of Ellipse, Given Foci: (-2,0) and (2,0) and Vertices: (-6,0) and (6,0).

Answer by stanbon(75887)

I have such a hard time doing this problem:
Write Standard Equation of Ellipse, Given Foci: (-2,0) and (2,0) and Vertices: (-6,0) and (6,0).
-----------------------
Draw the picture.
Do you see that a = 6 and c = 2.
For an ellipse, a^2 + b^2 = c^2, so b = sqrt[6^2-2^2] = sqrt(32) = 4sqrt(2)
----------------------------
EQUATION:
x^2/36 + y^2/32 = 1
=========================
Cheers,
Stan H.
=========================